### v2

@@ 3.2 Step Recognition
The step recognition module, denoted $\rho$, takes a window $W=\{{f_{{\tau}-w},..,f_{\tau}}\}$ with a fixed length $w$ of frames from $v$ to the current frame $\tau$ as input and is tasked with recognizing the last step performed $s^{\rho}_{\tau}$. We leverage MiniRoad [37] as $\rho$, which currently represents the state-of-the-art and is also computationally efficient in terms of GFlops and parameter count.
@@
During training, the loss for $\rho$ is calculated using a cross-entropy loss, comparing the actual step $s_{\tau}$ with the predicted one $s^{\rho}_{\tau}$ as detailed in [37].
@@
Despite being a state-of-the-art model, MiniRoad [37] performs suboptimally on procedural datasets, primarily due to two key factors: the shift in action distribution between the training and testing phases, and the inherent class imbalance within the dataset. The distribution shift occurs because the actions seen during training often differ from those encountered during testing, leading to poorer generalization. Additionally, Assembly101 contains a significant imbalance in class frequencies (see Figure 3 for Assembly101-O’s split), where common actions dominate the dataset while less frequent yet crucial procedural actions are underrepresented.
While MiniROAD [37] produces per-frame predictions, the anticipator expects a sequence of distinct actions as input. A standard approach for transforming per-frame predictions into action sequences involves aggregating identical neighboring predictions into a single action segment. However, MiniROAD’s performance issues introduce significant noise into the input sequence for the anticipator module, mainly due to inconsistencies in per-frame predictions, such as oscillating between similar actions, which leads to unstable action predictions and degrades results in subsequent stages.
@@
To address this, we propose a frame aggregation strategy that efficiently converts per-frame predictions into coherent action sequences, aiming to reduce noise and stabilize predictions. This results in a cleaner input for the anticipator module, dubbed $AG$ in Fig. 2.
In our experiments, we asses three strategies to perform aggregation. One approach (Non-Overlapping Mode Aggregation, NOMA) uses nonoverlapping windows, while the other two employ overlapping windows. These two overlapping methods differ based on which frame is selected for imputing the aggregation result: the last frame (Overlapping Mode Aggregation, OMA) or the central frame (Overlapping Central Mode Aggregation, OCMA) (See Figure 4). All approaches are designed to reduce noise and eliminate redundant consecutive predictions, ensuring that the resulting sequence aligns closely with the actual procedure measured by the Levenshtein Similarity. In section 5.2, we compare and discuss in details the aggregation strategies and find NOMA to be the most effective, which we therefore select.
@@ 3.3 Step Anticipation
We employ an LLM as our $\xi$ module for next-step prediction, using as input transcripts from procedural video. Our step anticipation operates without domain specific training or fine-tuning, leveraging only the reasoning from Automatic Chain of Thought (ACoT) [28] (see Fig. 2) and few-shot prompting method. These prompts consist of three components: the system prompt, the ACoT prompt, which enables reasoning, and the contextual transcripts $C$ together with the current aggregated sequence $AG$. $C$ is derived from similar transcripts (i.e., different videos for the object we are considering) to provide the LLM with information about typical step sequences and their order (see following examples). Sequence $S^{\rho}_{\tau}$, includes the current sequence of actions up to the current frame $f_{\tau-1}$, as they have been detected by our module $\rho$, i.e., $S^{\rho}_{\tau}=[s^{\rho}_{1},...,s^{\rho}_{\tau-1}]$, where $s^{\rho}_{i}$ represents the action detected by the module $\rho$ at frame $i$. $S^{\rho}_{\tau}$ then undergoes our aggregation strategy, which results in $AG$.
@@ 3.4 Mistake Detection
$\begin{cases}s^{\rho}_{\tau}\neq s^{\xi}_{\tau}&\text{MISTAKE}\\ s^{\rho}_{\tau}=s^{\xi}_{\tau}&\text{CORRECT}\\ \end{cases}$ (2)
@@ 4 Experiments
This section describes the benchmark datasets, adopted evaluation metrics and the main experimental results. We introduce Assembly101-O and Epic-tent-O as adaptations of the original datasets [2, 1], detailing the selected labeling criteria for online benchmarking, a novel configuration of training and test splits to accommodate open-set procedural mistakes. We define our proposed online metrics in Sec. 4.2. Finally, we introduce the main experimental results in Sec. 4.3, whereas further insights that motivate our design choices are provided through the ablation studies in Sec. 5.
@@ 4.1.1 Assembly101-O
Assembly101 [2] is a large-scale video dataset containing 362 videos of assembly and disassembly tasks involving 101 types of toy vehicles, recorded from eight multiple static and four egocentric cameras. It includes extensive annotations at various levels of granularity, with over 100K coarse action segments and 1M fine-grained action segments and 18M 3D hand poses. The dataset supports diverse challenges, including action anticipation and segmentation, mistake detection and 3D pose-based action recognition.
Assembly101 for Online and Open-Set Mistake Detection (Proposed) We introduce a novel split of the dataset [2] tailored explicitly for online, open-set mistake detection. Assembly101-O introduces two main modifications to [2] as we show in Figure 5: a new train/test split and revised procedure lengths.
@@ 4.1.2 Epic-tent-O
Epic-tent for Online and Open-Set Mistake Detection (Proposed) We propose a novel split for the Epic-tent dataset [1] tailored for open-set mistake detection. The dataset includes annotations for nine distinct mistake types, though only four of these categories —“order”, “omit”, “correction” and “repeat” — are considered procedural mistakes, as they represent deviations that compromise the task’s procedural integrity. Other categories such as “slow”, “search”, “misuse”, “motor” and “failure” are not procedural errors, as they do not disrupt the task’s progression.
Unlike Assembly101 [2], Epic-tent is designed explicitly for supervised error detection with each recorded procedure containing at least one mistake, making it incompatible with the split procedure proposed for Assembly101-O. However, Epic-tent provides confidence scores from each participant, reflecting their self-assessed uncertainty during task execution. Therefore, we define a custom split strategy where videos from higher-confidence participants form the training set. In contrast, participants showing more significant uncertainty, and thus potentially prone to more errors, make up the test set. In practice, the videos that form the test set have a median confidence score under 0.6, while the others form the train set. Only 22 videos have the confidence score annotations, while the remaining 7 do not and are assigned to the test set. This partition results in 14 training videos and 15 test videos.
@@ 4.2 Evaluation Definition
To mitigate this imbalance, we utilize Balanced Precision [23], denoted $\mathrm{Prec}_{b}$, by scaling the contribution of each False Positive by the ratio of positive-to-negative samples. Concretely, if $N_{+}$ is the total number of positive (error) samples and $N_{-}$ is the total number of negative (correct) samples, we define:
@@
[…]
@@ 4.3.1 Compared Methods
OadTR for mistake detection The work in [14] presents a framework for online action detection called OadTR, which uses a Vision Transformer to capture video clips’ temporal structure and context. The framework consists of an encoder-decoder architecture: the encoder processes historical observations to output a task token that represents the current action, while the decoder refines this task token by incorporating information from anticipated future clips. In the context of procedural error detection, we consider a mistake to have occurred if the output from the encoder does not match the output from the decoder, indicating a divergence between observed and expected actions.
BERT-based Mistake Detection We also leverage the capabilities of BERT [56], using its specific [CLS] token to predict whether an action sequence is correct or erroneous. More specifically, we fine-tune BERT with a next-token-prediction task, training the model to determine whether step B logically follows step A within a given procedure. In our setting, each step is represented as a set of two words, such as attach wheel, which describes coarse-level actions. During evaluation, BERT is presented with pairs of actions and predicts whether the sequence of those actions is correct. The rationale of BERT lies in its pre-training on a vast corpus of text, which helps it understand contextual relationships between procedural steps.
Learned Task Graph Representations We also evaluate an approach based on explicit task graph representations [16]. Procedural activities consist of sequences of key steps that work toward specific goals, making task graphs a valuable, human-understandable representation for capturing the partial ordering of these steps. Unlike traditional methods that rely on hand-crafted procedures to extract task graphs from videos, this approach uses maximum likelihood optimization of edge weights to directly learn task graphs, allowing gradient-based learning that can integrate seamlessly with neural networks. While effective, this baseline is not directly comparable to our method, as it requires access to the entire dataset during training to extract the task graph, whereas our approach relies only on few-shot inference for anticipating actions. We refer to the “Direct Optimization” method trained with the “Task-Graph Maximul Likelihood” loss proposed in [16] as TGML-DO.
@@ 4.3.2 Results
[…]
The TGML-DO method [16], which learns task graph representations through gradient-based optimization, also demonstrates strong performance. However, unlike our anticipator, which operates in a few-shot inference setting, TGML-DO checks whether the observed action is correct by referencing domain-specific procedural knowledge mined from the entire dataset during training. This distinction likely contributes to the superior generalization of our model, particularly when using predicted steps instead of oracle labels in the Assembly101-O dataset. Our method proves to be more robust to procedural variations and better suited for handling previously unseen cases.
[…]
@@
BERT [56] Oracle 93.2 20.0 32.9 71.6 5.6 10.4
PREGO Oracle GPT-3.5 51.4 87.5 64.7 25.6 90.6 39.9
PREGO [22] Oracle DeepSeek-R1 51.7 98.3 67.8 23.8 100 38.4
@@
TGML-DO [16] Oracle 77.82 90.4 83.64 91.24 27.3 42.03
TGML-DO [16] MiniRoad [37] 53.77 37.3 44.05 96.55 14.1 24.60
OadTR for MD [14] OadTR [14] OadTR [14] 50.9 18.1 26.7 40.4 21.7 28.3
PREGO MiniRoad [37] GPT-3.5 50.2 75.8 61.4 19.5 73.3 29.4
PREGO [22] MiniRoad [37] DeepSeek-R1 51.2 96.5 65.7 20.3 93.3 33.6
TI-PREGO MiniRoad [37] Llama 3.1 51.5 97.2 67.4 21.2 93.3 34.6
@@ 5.1.1 Model Selection
[…]

### v3

@@ 3.2 Step Recognition
The step recognition module, denoted $\rho$, takes a window $W=\{{f_{{\tau}-w},..,f_{\tau}}\}$ with a fixed length $w$ of frames from $v$ to the current frame $\tau$ as input and is tasked with recognizing the last step performed $s^{\rho}_{\tau}$. We leverage MiniROAD [An2023MiniROADMR] as $\rho$, which currently represents the state-of-the-art and is also computationally efficient in terms of GFlops and parameter count.
@@
During training, the loss for $\rho$ is calculated using a cross-entropy loss, comparing the actual step $s_{\tau}$ with the predicted one $s^{\rho}_{\tau}$ as detailed in [An2023MiniROADMR].
@@
Despite being a state-of-the-art model, MiniROAD [An2023MiniROADMR] performs suboptimally on procedural datasets, primarily due to two key factors: the shift in action distribution between the training and testing phases, and the inherent class imbalance within the dataset. The distribution shift occurs because the actions seen during training often differ from those encountered during testing, leading to poorer generalization. Additionally, Assembly101 contains a significant imbalance in class frequencies (see Figure 3 for Assembly101-O’s split), where common actions dominate the dataset while less frequent yet crucial procedural actions are underrepresented.
While MiniROAD [An2023MiniROADMR] produces per-frame predictions, the anticipator expects a sequence of distinct actions as input. A standard approach for transforming per-frame predictions into action sequences involves aggregating identical neighboring predictions into a single action segment. However, MiniROAD’s performance issues introduce significant noise into the input sequence for the anticipator module, mainly due to inconsistencies in per-frame predictions, such as oscillating between similar actions, which leads to unstable action predictions and degrades results in subsequent stages.
@@
To address this, we propose three frame aggregation strategies that efficiently converts per-frame predictions into coherent action sequences, aiming to reduce noise and stabilize predictions. This results in a cleaner input for the anticipator module, dubbed $AG$ in Fig. 2.
Non-Overlapping Mode Aggregation (NOMA)
The first aggregation strategy we propose (Figure 4) consists of a two-step process. In the first step, we calculate the mode for each non-overlapping sliding window of frames of a given length $L$. Since the windows are non-overlapping, this approach introduces a slight delay in processing, which is a trade-off for improved prediction stability. The mode, representing the most frequent action label within that window, replaces all predictions in that window, ensuring consistency across the frames and removing the noise of smaller portions wrongly predicted. In the second step, we apply the elimination of successive duplicates to remove consecutive repeated action labels, which prevents redundant predictions. This cleaned output serves as the input to the anticipation module, ultimately aiming to produce a more stable and accurate sequence of action predictions.
Overlapping Mode Aggregation (OMA)
The second aggregation strategy we propose (Fig. 4) employs a sliding window approach with a stride of $1$ frame, allowing for overlapping portions of frames. In this method, we evaluate the mode within the overlapping window that includes the last frame and the previous L1L-1 frames. The prediction of the last frame is then substituted with this mode, which captures the most frequent action label in that localized area. This method is designed with the intent to reduce delay, as it processes frames more dynamically compared to non-overlapping windows.
Overlapping Centered Mode Aggregation (OCMA)
NOMA uses nonoverlapping windows, while the other two employ overlapping windows. These two overlapping methods differ based on which frame is selected for imputing the aggregation result: the last frame (Overlapping Mode Aggregation, OMA) or the central frame (Overlapping Central Mode Aggregation, OCMA) (See Figure 4). All approaches are designed to reduce noise and eliminate redundant consecutive predictions, ensuring that the resulting sequence aligns closely with the actual procedure measured by the Levenshtein Similarity.
In Sec. 5.2, we assess, compare and discuss in details the impact of each aggregation strategy and find NOMA to be the most effective, which we therefore select for our proposal.
@@ 3.3 Step Anticipation
We employ an LLM as our $\xi$ module for next-step prediction, using as input transcripts from procedural video. Our step anticipation operates without domain specific training or fine-tuning, leveraging only the reasoning from Automatic Chain of Thought (ACoT) [Zhang2022AutomaticCO] (see Fig. 2) and few-shot prompting method. These prompts consist of three components: the system prompt, the ACoT prompt, which enables reasoning, and the contextual transcripts $C$ together with the current aggregated sequence $AG$. $C$ is derived from similar transcripts (i.e., different videos for the object we are considering) to provide the LLM with information about typical step sequences and their order (see following examples). Sequence $S^{\rho}_{\tau}$, includes the current sequence of actions up to the current frame $f_{\tau-1}$, as they have been detected by our module $\rho$, i.e., $S^{\rho}_{\tau}=[s^{\rho}_{1},...,s^{\rho}_{\tau-1}]$, where $s^{\rho}_{i}$ represents the action detected by the module $\rho$ at frame $i$. $S^{\rho}_{\tau}$ then undergoes our aggregation strategy, which results in $AG$.
@@ 3.4 Mistake Detection
[…]
@@ 4 Experiments
This section describes the benchmark datasets, adopted evaluation metrics and the main experimental results. We introduce Assembly101-O and Epic-tent-O as adaptations of the original datasets [sener2022assembly101, epicTent], detailing the selected labeling criteria for online benchmarking, a novel configuration of training and test splits to accommodate open-set procedural mistakes. We define our proposed online metrics in Sec. 4.2. Finally, we introduce the main experimental results in Sec. 4.3, whereas further insights that motivate our design choices are provided through the ablation studies in Sec. 5.
@@ 4.1.1 Assembly101-O
Assembly101 [sener2022assembly101] is a large-scale video dataset containing 362 videos of assembly and disassembly tasks involving 101 types of toy vehicles, recorded from eight multiple static and four egocentric cameras. It includes extensive annotations at various levels of granularity, with over 100K coarse action segments and 1M fine-grained action segments and 18M 3D hand poses. The dataset supports diverse challenges, including action anticipation and segmentation, mistake detection and 3D pose-based action recognition.
Assembly101 for Online and Open-Set Mistake Detection (Proposed) We introduce a novel split of the dataset [sener2022assembly101] tailored explicitly for online, open-set mistake detection. Assembly101-O introduces two main modifications to [sener2022assembly101] as we show in Figure 5: a new train/test split and revised procedure lengths.
@@ 4.1.2 Epic-tent-O
Epic-tent for Online and Open-Set Mistake Detection (Proposed) We propose a novel split for the Epic-tent dataset [epicTent] tailored for open-set mistake detection. The dataset includes annotations for nine distinct mistake types, though only four of these categories —“order”, “omit”, “correction” and “repeat” — are considered procedural mistakes, as they represent deviations that compromise the task’s procedural integrity. Other categories such as “slow”, “search”, “misuse”, “motor” and “failure” are not procedural errors, as they do not disrupt the task’s progression.
Unlike Assembly101 [sener2022assembly101], Epic-tent is designed explicitly for supervised error detection with each recorded procedure containing at least one mistake, making it incompatible with the split procedure proposed for Assembly101-O. However, Epic-tent provides confidence scores from each participant, reflecting their self-assessed uncertainty during task execution. Therefore, we define a custom split strategy where videos from higher-confidence participants form the training set. In contrast, participants showing more significant uncertainty, and thus potentially prone to more errors, make up the test set. In practice, the videos that form the test set have a median confidence score under 0.6, while the others form the train set. Only 22 videos have the confidence score annotations, while the remaining 7 do not and are assigned to the test set. This partition results in 14 training videos and 15 test videos.
@@ 4.2 Evaluation Definition
To mitigate this imbalance, we utilize Balanced Precision [Wang2021OadTROA], denoted $\mathrm{Prec}_{b}$, by scaling the contribution of each False Positive by the ratio of positive-to-negative samples. Concretely, if $N_{+}$ is the total number of positive (error) samples and $N_{-}$ is the total number of negative (correct) samples, we define:
@@
Finally, we define a Balanced F1 Score, denoted $\text{F1 score}_{b}$, which is the harmonic mean of $\mathrm{Prec}_{b}$ and $\mathrm{Recall}$:
$\text{F1 score}_{b}\;=\;2\,\frac{\mathrm{Prec}_{b}\times\mathrm{Recall}}{\mathrm{Prec}_{b}+\mathrm{Recall}}\,.$ (5)
These balanced metrics specifically assess the model’s ability to distinguish correct from incorrect steps, without being influenced by heavily skewed settings. By adjusting the impact of false positives based on the imbalance between positive and negative classes, $\mathrm{Prec}_{b}$ and $\text{F1 score}_{b}$ provide a fairer assessment of mistake detection when the negative class vastly outnumbers the positive class.
@@ 4.3.1 Compared Methods
OadTR for mistake detection The work in [OadTR] presents a framework for online action detection called OadTR, which uses a Vision Transformer to capture video clips’ temporal structure and context. The framework consists of an encoder-decoder architecture: the encoder processes historical observations to output a task token that represents the current action, while the decoder refines this task token by incorporating information from anticipated future clips. In the context of procedural error detection, we consider a mistake to have occurred if the output from the encoder does not match the output from the decoder, indicating a divergence between observed and expected actions.
BERT-based Mistake Detection We also leverage the capabilities of BERT [Devlin2019BERTPO], using its specific [CLS] token to predict whether an action sequence is correct or erroneous. More specifically, we fine-tune BERT with a next-token-prediction task, training the model to determine whether step B logically follows step A within a given procedure. In our setting, each step is represented as a set of two words, such as attach wheel, which describes coarse-level actions. During evaluation, BERT is presented with pairs of actions and predicts whether the sequence of those actions is correct. The rationale of BERT lies in its pre-training on a vast corpus of text, which helps it understand contextual relationships between procedural steps.
Learned Task Graph Representations We also evaluate an approach based on explicit task graph representations [seminara2024differentiable]. Procedural activities consist of sequences of key steps that work toward specific goals, making task graphs a valuable, human-understandable representation for capturing the partial ordering of these steps. Unlike traditional methods that rely on hand-crafted procedures to extract task graphs from videos, this approach uses maximum likelihood optimization of edge weights to directly learn task graphs, allowing gradient-based learning that can integrate seamlessly with neural networks. While effective, this baseline is not directly comparable to our method, as it requires access to the entire dataset during training to extract the task graph, whereas our approach relies only on few-shot inference for anticipating actions. We refer to the “Direct Optimization” method trained with the “Task-Graph Maximul Likelihood” loss proposed in [seminara2024differentiable] as TGML-DO.
@@ 4.3.2 Results
Table 2 presents the main results in terms of $\mathrm{Precision}_{b}$, $\mathrm{Recall}$, and $\mathrm{\text{F1 score}}_{b}$.
In the upper section of the table, we report results obtained using an oracle step-recognizer, which serves as an upper bound. Here, the recognition module perfectly aligns with the ground truth, ensuring that errors stem solely from the anticipation module. The One-step memory method performs poorly as it relies only on the preceding action. Even if BERT adopts a more abstract reasoning strategy, it suffers from a conservative bias that leads to very low recall, ultimately resulting in a significantly lower $\mathrm{\text{F1 score}}_{b}$. PREGO, on the other hand, incorporates symbolic reasoning to capture richer contextual information, with TI-PREGOLlama achieving the highest performance across most metrics. To ensure a fair comparison, we also evaluated the baseline PREGO model with Llama 3.1. As shown in Table 2, TI-PREGO (74.7% $\mathrm{\text{F1 score}}_{b}$) outperforms the PREGO baseline (67.5% $\mathrm{\text{F1 score}}_{b}$) on Assembly101-O, confirming that our methodological improvements are the primary driver of the performance gain.
The TGML-DO method [seminara2024differentiable], which learns task graph representations through gradient-based optimization, also demonstrates strong performance. However, unlike our anticipator, which operates in a few-shot inference setting, TGML-DO checks whether the observed action is correct by referencing domain-specific procedural knowledge mined from the entire dataset during training. This distinction likely contributes to the superior generalization of our model, particularly when using predicted steps instead of oracle labels in the Assembly101-O dataset. Our method proves to be more robust to procedural variations and better suited for handling previously unseen cases.
In the lower section of the table, we evaluate models using state-of-the-art step recognizers instead of an oracle. The OadTR model, designed for offline mistake detection, shows limited performance due to its reliance on fixed video segmentation, resulting in a low $\mathrm{\text{F1 score}}_{b}$ of 26.7% on Assembly101-O and 28.3% on Epic-tent-O. By contrast, combining MiniRoad with large language models for step anticipation leads to consistent gains. When paired with GPT-3.5, PREGO achieves a 34,7% absolute improvement in $\mathrm{\text{F1 score}}_{b}$ over OadTR on Assembly101-O. The use of DeepSeek-R1 yields similar gains. The best performance is obtained by TI-PREGO combined with Llama 3.1, which achieves an $\mathrm{\text{F1 score}}_{b}$ of 74.7% on Assembly101-O and 57.8% on Epic-tent-O, outperforming all other configurations. This direct comparison, using the same Llama 3.1 model, confirms our method’s advantage: TI-PREGO (67.4% $\mathrm{\text{F1 score}}_{b}$) outperforms the PREGO baseline (67.1% $\mathrm{\text{F1 score}}_{b}$) on Assembly101-O, with a similar trend on Epic-tent-O (34.6% vs. 33.7%). This highlights the robustness of Llama 3.1 in modeling procedural structure and anticipating upcoming steps, particularly in open-set, real-world mistake detection scenarios.
@@
BERT [Devlin2019BERTPO] Oracle 93.2 20.0 32.9 71.6 5.6 10.4
PREGO [Flaborea_2024_CVPR] Oracle GPT-3.5 51.4 87.5 64.7 25.6 90.6 39.9
PREGO [Flaborea_2024_CVPR] Oracle DeepSeek-R1 51.7 98.3 67.8 23.8 100 38.4
PREGO [Flaborea_2024_CVPR] Oracle Llama 3.1 56.8 89.6 69.5 28.0 92.0 42.9
@@
TGML-DO [seminara2024differentiable] Oracle 77.82 90.4 83.64 91.24 27.3 42.03
TGML-DO [seminara2024differentiable] MiniRoad [An2023MiniROADMR] 53.77 37.3 44.05 96.55 14.1 24.60
OadTR for MD [OadTR] OadTR [OadTR] OadTR [OadTR] 50.9 18.1 26.7 40.4 21.7 28.3
PREGO [Flaborea_2024_CVPR] MiniRoad [An2023MiniROADMR] GPT-3.5 50.2 75.8 61.4 19.5 73.3 29.4
PREGO [Flaborea_2024_CVPR] MiniRoad [An2023MiniROADMR] DeepSeek-R1 51.2 96.5 65.7 20.3 93.3 33.6
PREGO [Flaborea_2024_CVPR] MiniRoad [An2023MiniROADMR] Llama 3.1 51.3 96.8 67.1 20.6 93.3 33.7
TI-PREGO MiniRoad [An2023MiniROADMR] Llama 3.1 51.5 97.2 67.4 21.2 93.3 34.6
@@ 5.1.1 Model Selection
In this analysis, we evaluate the performance of various large language models (LLMs), as shown in Table 3, using a few-shot context. The results reveal a clear performance hierarchy among the tested models. In particular, LLAMA 3.1 8B, an open-source model less computationally expensive during inference, demonstrates robust performance (see Table 2). LLAMA 3.1 8B achieved the highest $\mathrm{\text{Precision}}_{b}$ at 56.70% and the best $\mathrm{\text{F1 score}}_{b}$of 69.4%, significantly outperforming its counterparts.